OpenStudy (anonymous):

Hi! Can anyone help me with these questions? Please and thank you! 3. A horse travels 40 kilometers per hour faster than a mule. The horse goes 360 kilometers in three hours less time than the mule goes 360 kilometers. Find the speed of each animal? 4. An airplane flies 900 miles against a headwind of 25 miles per hour. The plane took 15minutes longer for this flight than with a tailwind of 25 miles per hour. How fast could the plane fly in still air?

OpenStudy (anonymous):

for problem (3), let u=speed of horse in km/h and v=speed of mule in km/h, from 1st condition, u-v=40 .............................(a) from second condition, 360/v-360/u=3 =>1/v-1/u=3/360=1/120 .......................(b) solve (a) and (b),

OpenStudy (anonymous):

for problem (4), let u=speed of airplane in still air in miles/h, v= speed of wind in miles/h from 1st condition, u-v=25 ........................(a) from 2nd condition, 900/(u-v)-900/(u+v)=15/60 => 900/25-900/(u+v)=1/4 => 1/(u+v)=1/25-1/4(900) => u+v=143/25 => u+u-25=143/25 => 2u=143/25+25=768/25 => u=768/25(2)=384/25=15.36 so u=15.36 miles/h

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome